OpenStudy (anonymous):

Slope of -3 and through (8,5)

OpenStudy (anonymous):

@Jsjtw19 so whats the q/n?r u talking abt eqn?

OpenStudy (anonymous):

y-5=-3(x-8) y-5=-3x+24 y+3x=29 is the answer!

OpenStudy (anonymous):

trying to find the equation of a line?

OpenStudy (anonymous):

isn't tht correct @Peter14 ?

OpenStudy (anonymous):

if you need it in slope intercept form, it is y = -3x + 29. or if you need it in standard form, it is 3x + y = 29

OpenStudy (anonymous):

as simple as that!

OpenStudy (anonymous):

nice job ayumi :)

OpenStudy (anonymous):

yes, that's right

OpenStudy (anonymous):

:)) @kelliegirl33

OpenStudy (anonymous):

:D @Peter14

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

thats what we are here for :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

ur welcome@Jsjtw!

OpenStudy (anonymous):

@Justine1998 whts d q/n?